OpenStudy (anonymous):

The perimeter of a triangle is 51 meters. the second is twice as long as the first side. the third side is 3 meters less then three times the first What are the lengths of the sides of the triangle? (Thanks)

OpenStudy (anonymous):

a+b+c=51 b=2a c=3a-3 a+2a+3a-3=51 6a-3=51 6a=54 a=9; b=18 c=24